OpenStudy (anonymous):

Use inverse trigonometric functions to find the solutions of the equation that are in the given interval, and approximate the solutions to four decimal places. cos^2(x)+2cosx-2=0 [0, 2pi] I'm having a little trouble with this. I believe I'm supposed to solve for cosine of x and then use the inverse cosine function to find x in radians?

OpenStudy (anonymous):

So you have: \[\cos^2(x)+2\cos(x)-2=0\] This is actually a quadratic in disguise! Let \([u=\cos(x)]\). You have: \[\eqalign{ &u^2+2u-2=0 \\ }\] So then you can solve for \(u\) using the quad formula: \[u=\frac{-2\pm\sqrt{4+4}}{2}=\frac{-2\pm\sqrt{8}}{2}=\frac{-2\pm2\sqrt{2}}{2}=-1\pm\sqrt{2}\] So therefore, you can back substitute: \[\cos(x)=-1\pm\sqrt{2}\] Now, before I go any further, what grade is this for, or rather which class?

OpenStudy (anonymous):

Its a trigonometry and analytic geometry course

OpenStudy (anonymous):

In which grade?

OpenStudy (anonymous):

Its a freshman college course

OpenStudy (anonymous):

Its the same as a high school trig and precalc course essentially

OpenStudy (anonymous):

Then you need to realize the intervals of \(\theta\) where \(\cos(x)\) is positive and negative:  So then you know that if \(\cos(x)=+ve\), then it must be in the first or last quadrant. Also, if \(\cos(x)=-ve\), then it must be in the second or third quadrant.